# Safe But Not Sorry: Reducing Over-Conservatism in Safety Critics via Uncertainty-Aware Modulation

### v1

@@ 4.2. Uncertainty Refinement
Method
Environment DDPG Safety Critic Conservative Safety Critic Uncertain Safety Critic (NR) Uncertain Safety Critic
Safety Gymnasium Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$   Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
CarGoal1 ($\chi=1.0$) $6.57\pm 0.45$ $1.05\pm 0.39$ $6.60\pm 0.45$ $1.02\pm 0.37$ $6.56\pm 0.40$ $1.03\pm 0.39$ $6.56\pm 0.39$ $1.06\pm 0.41$ $6.55\pm 0.39$ $0.96\pm 0.33$
CarGoal2 ($\chi=5.0$) $8.74\pm 0.72$ $5.36\pm 1.47$ $7.54\pm 1.38$ $5.65\pm 2.00$ $8.65\pm 0.75$ $5.28\pm 1.44$ $8.84\pm 0.64$ $5.44\pm 1.43$ $8.84\pm 0.62$ $5.05\pm 1.27$
CarButton1 ($\chi=5.0$) $6.20\pm 0.74$ $9.50\pm 3.01$ $5.12\pm 1.89$ $5.96\pm 2.68$ $48\pm 1.83 5.5$$57\pm 2.63$ $6.31\pm 0.55$ $6.37\pm 2.05$ $6.38\pm 0.62$ $6.01\pm 2.13$
CarButton2 ($\chi=5.0$) $7.51\pm 0.67$ $10.30\pm 3.32$ $2.48\pm 3.32$ $4.04\pm 4.46$ $6.45\pm 2.44$ $6.49\pm 3.19$ $7.49\pm 0.79$ $8.02\pm 2.52$ $7.69\pm 0.62$ $6.32\pm 2.23$
Gymnasium Robotics Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$   Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
FetchReach ($\chi=5.0$) $4.36\pm 16.04$ $8.19\pm 7.70$ $1.98\pm 4.65$ $3.18\pm 1.05$ $1.53\pm 6.51$ $3.91\pm 3.07$ $0.55\pm 4.00$ $3.34\pm 1.36$ $3.92\pm 2.51$ $3.20\pm 1.11$
Mujoco Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$   Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
HalfCheetah ($\chi=0.1$) $5.71\pm 2.48$ $21.82\pm 23.20$ $3.19\pm 3.43$ $0.32\pm 0.63$ $4.13\pm 1.29$ $2.71\pm 5.70$ $6.64\pm 2.26$ $0.62\pm 0.51$ $3.76\pm 2.70$ $0.20\pm 0.40$
@@ 5. Evaluation
We evaluate USC  [^1] Code available at: https://anonymous.4open.science/r/USC-65DB/README.md using tasks from the Safety Gymnasium ji2023safety (19), Gymnasium Robotics gymrobot (26), and Mujoco towers2024gymnasium (38) test-suites. These benchmarks assess performance and safety in high-dimensional scenarios with various constraints, enabling a fair comparison of USC against approaches also used similar benchmarks bharadhwaj2020conservative (12). The selected tasks cover varying complexities in constraints, goals, and agent types:
  • CarGoal1: A car-like wheeled agent must reach a goal while avoiding nine hazards, all with randomised positions.
  • CarGoal2: Similar to the CarGoal1 environment, but with twenty hazards.
  • CarButton1: An agent must press the correct button while avoiding eight hazards and three incorrect buttons, all with randomised positions.
  • CarButton2: Similar to the CarButton1 environment, but with fourteen hazards.
@@
  • DDPG: A deep deterministic policy gradient agent ddpg-paper (23) which is the foundational baseline.
  • Safety Critic (CSC): A DDPG agent leveraging a safety critic to evaluate safety srinivasan2020learning (34).
  • Conservative Safety Critic (CSC): A DDPG agent with a conservative safety critic that uses conservative estimates to evaluate safety bharadhwaj2020conservative (12).
We also ablate USC by comparing against Uncertain Safety Critic (NR), a USC variant without the uncertainty refinement stage. Other safe exploration approaches discussed in Section 2 were excluded because they represent earlier research from approaches used in our evaluation. To enable a fair comparison, all methods use the same DDPG configuration; algorithm details and hyperparameter information are provided in Appendix D. Following bharadhwaj2020conservative (12), we perform five independent runs, with results showing mean scores and confidence intervals (standard error).
@@ 5.1. Performance Results
Method
Environment DDPG Safety Critic Conservative Safety Critic Uncertain Safety Critic (NR) Uncertain Safety Critic
Safety Gymnasium Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
CarGoal1 ($\chi=1.0$) $7.90\pm 2.40$ $3.10\pm 4.11$ $7.57\pm 1.89$ $4.80\pm 5.93$ $7.21\pm 2.64$ $2.00\pm 6.00$ $5.79\pm 2.34$ $2.60\pm 5.66$ $7.19\pm 1.97$ $0.70\pm 2.10$
CarGoal2 ($\chi=5.0$) $9.13\pm 2.88$ $41.80\pm 40.59$ $11.48\pm 2.51$ $31.20\pm 33.28$ $9.44\pm 2.79$ $27.00\pm 17.49$ $9.38\pm 3.95$ $26.30\pm 26.71$ $9.48\pm 3.30$ $20.20\pm 24.63$
CarButton1 ($\chi=5.0$) $5.36\pm 1.87$ $71.40\pm 126.01$ $2.63\pm 2.14$ $0.00\pm 0.00$ $3.96\pm 1.62$ $7.90\pm 15.93$ $7.65\pm 2.66$ $22.20\pm 25.96$ $6.41\pm 2.53$ $15.90\pm 28.40$
CarButton2 ($\chi=5.0$) $9.09\pm 2.69$ $31.30\pm 24.58$ $11\pm 0.17 0.0$$00\pm 0.00$ $6.33\pm 2.95$ $14.10\pm 17.39$ $9.59\pm 2.59$ $24.00\pm 41.23$ $8.51\pm 4.79$ $13.60\pm 13.71$
Gymnasium Robotics Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
FetchReach ($\chi=5.0$) $4.74\pm 0.10$ $3.09\pm 1.22$ $4.81\pm 0.09$ $3.12\pm 1.09$ $4.78\pm 0.10$ $2.98\pm 1.17$ $4.78\pm 0.13$ $2.97\pm 1.75$ $4.78\pm 0.07$ $2.45\pm 1.25$
Mujoco Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
HalfCheetah ($\chi=0.1$) $9.24\pm 2.17$ $69.80\pm 27.32$ $6.91\pm 0.47$ $0.50\pm 0.92$ $8.66\pm 1.57$ $16.00\pm 10.11$ $5.74\pm 0.28$ $0.60\pm 0.49$ $5.44\pm 0.42$ $0.20\pm 0.40$
@@
Method
Environment DDPG SC CSC USC (NR) USC
CarGoal1 ($\chi=1.0$) $\% 100100\%$ $\% 100100\%$ $100\%$
CarGoal2 ($\chi=5.0$) $\% 100100\%$ $\% 100100\%$ $100\%$
CarButton1 ($\chi=5.0$) $90\%$ $50\%$ $70\%$ $\% 100100\%$
CarButton2 ($\chi=5.0$) $100\%$ $0\%$ $80\%$ $\% 100100\%$
FetchReach ($\chi=5.0$) $\% 100100\%$ $\% 100100\%$ $100\%$
@@ B.1. Theorem 1
Theorem 1. The log-exp-sum (LSE) penalty, used in Equation 3, is more effective at encouraging conservatism when the $m$ alternative actions are sampled uniformly over the action space, rather than from the current policy bharadhwaj2020conservative (12). If the policy aims to avoid high-cost regions, then the uniform sampling is more likely to draw from high-cost out-of-distribution actions. Formally, for any fixed state $s$, the conservative term is strictly larger under uniform scaling.
@@ C.1. Extended Analysis of Core Results
The deployment-phase evaluations, see in rFigure 9, confirm the trends reported in Table 3. Across all CarGoal and CarButton tasks, USC consistently delivers higher reward while keeping costs relatively low compared to all other methods, and it maintains the highest average distance to hazards. Safety Critic again proves brittle, with frequent cost spikes and poor hazard avoidance, while CSC is more stable but overly conservative, often suppressing reward. Notably, in the more difficult CarButton tasks, USC is the only safety-aware method that sustains balanced performance without collapsing into unsafe or overly cautious behaviour and achieves 100% success rate, whereas the other safety-aware methods fail to reach the goal every time. These results reinforce that USC’s uncertainty-modulated training advantage translates into reliable and safe deployment performance.
@@ Appendix D Hyperparameter Analysis and Computational Overheads
Hyperparameters for the DDPG algorithm started with author recommendations ddpg-paper (23). They were manually tuned afterwards to achieve a high performance on individual environments before tests were carried out, meaning the RL algorithm for all approaches was of high performance and fair comparison.

### v2

@@ 4.2. Uncertainty Refinement
Safety Gymnasium Gymnasium Robotics Mujoco
Technique Metric Goal1 ($\chi=1.0$) Goal2 ($\chi=5.0$) Button1 ($\chi=5.0$) Button2 ($\chi=5.0$) FetchReach ($\chi=5.0$) HalfCheetah ($\chi=0.1$)
DDPG Reward $\uparrow$ $6.57\pm 0.47$ $8.74\pm 0.72$ $6.20\pm 0.74$ $7.51\pm 0.67$ $4.36\pm 16.04$ $5.71\pm 2.48$
Cost $\downarrow$ $1.05\pm 0.39$ $5.36\pm 1.47$ $9.50\pm 3.01$ $10.30\pm 3.32$ $8.19\pm 7.70$ $21.82\pm 23.20$
PID-Lag Reward $\uparrow$ $6.48\pm 0.49$ $5.88\pm 3.39$ $2.49\pm 2.57$ $1.30\pm 2.61$ $1.43\pm 5.96$ $4.69\pm 1.52$
Cost $\downarrow$ $1.14\pm 0.51$ $4.94\pm 2.27$ $4.50\pm 4.26$ $3.40\pm 4.77$ $3.45\pm 3.12$ $0.36\pm 0.35$
CPO Reward $\uparrow$ $6.40\pm 0.76$ $8.74\pm 0.81$ $5.99\pm 1.20$ $7.56\pm 0.79$ $1.80\pm 4.49$ $3.69\pm 1.89$
Cost $\downarrow$ $1.21\pm 0.51$ $5.08\pm 1.26$ $7.47\pm 2.62$ $9.05\pm 2.80$ $4.24\pm 1.70$ $1.71\pm 1.93$
RCPO Reward $\uparrow$ $6.56\pm 0.38$ $8.28\pm 0.95$ $4.21\pm 2.38$ $3.67\pm 3.18$ $0.77\pm 6.57$ $6.28\pm 2.13$
Cost $\downarrow$ $1.02\pm 0.40$ $5.13\pm 1.71$ $5.46\pm 3.35$ $4.58\pm 3.72$ $3.84\pm 2.36$ $0.94\pm 0.84$
SC Reward $\uparrow$ $6.60\pm 0.45$ $7.54\pm 1.38$ $5.12\pm 1.89$ $2.48\pm 3.32$ $1.98\pm 4.65$ $3.19\pm 3.43$
Cost $\downarrow$ $1.02\pm 0.37$ $5.65\pm 2.00$ $5.96\pm 2.68$ $4.04\pm 4.46$ $3.18\pm 1.05$ $0.32\pm 0.63$
CSC Reward $\uparrow$ $6.56\pm 0.40$ $8.65\pm 0.75$ $5.48\pm 1.83$ $6.45\pm 2.44$ $1.53\pm 6.51$ $4.13\pm 1.29$
Cost $\downarrow$ $1.03\pm 0.39$ $28\pm 1.44 5.5$$57\pm 2.63$ $6.49\pm 3.19$ $3.91\pm 3.07$ $2.71\pm 5.70$
USC (NR) Reward $\uparrow$ $6.56\pm 0.39$ $8.74\pm 0.78$ $6.31\pm 0.55$ $7.49\pm 0.79$ $0.55\pm 4.00$ $6.64\pm 2.26$
Cost $\downarrow$ $1.06\pm 0.41$ $5.53\pm 1.43$ $6.37\pm 2.05$ $8.02\pm 2.52$ $3.34\pm 1.36$ $0.62\pm 0.51$
USC Reward $\uparrow$ $6.55\pm 0.39$ $8.84\pm 0.62$ $6.38\pm 0.62$ $7.69\pm 0.62$ $3.92\pm 2.51$ $3.76\pm 2.70$
Cost $\downarrow$ $0.96\pm 0.33$ $5.05\pm 1.27$ $6.02\pm 2.13$ $6.32\pm 2.23$ $3.20\pm 1.11$ $0.20\pm 0.40$
@@ 5. Evaluation
We evaluate USC  [^1] Code available at: https://github.com/team-daniel/USC using tasks from the Safety Gymnasium ji2023safety, Gymnasium Robotics gymrobot, and Mujoco towers2024gymnasium test-suites. These assess performance and safety in high-dimensional scenarios with various constraints, enabling a fair comparison of USC against approaches that also use similar benchmarks bharadhwaj2020conservative. The selected tasks cover varying complexities in constraints, goals, and agent types:
  • CarGoal1: A car-like wheeled agent must reach a goal while avoiding 9 hazards, all with randomised positions.
  • CarGoal2: Similar to CarGoal1, but with 20 hazards.
  • CarButton1: An agent must press the correct button, avoiding 8 hazards and 3 incorrect buttons, all with randomised positions.
  • CarButton2: Similar to the CarButton1, with 14 hazards.
@@
  • DDPG: A deep deterministic policy gradient agent ddpg-paper which is the foundational baseline.
  • PID-Lag: A Lagrangian constraint penalty using PID control to stabilise training stooke2020responsive.
  • CPO: A trust-region method that enforces constraints during each policy update achiam2017constrained.
  • RCPO: A method which incorporates constraint costs into the reward via a learned Lagrange multiplier tessler2018reward.
  • Safety Critic (CSC): A DDPG agent leveraging a safety critic to evaluate safety srinivasan2020learning.
  • Conservative Safety Critic (CSC): A DDPG agent with a conservative safety critic that uses conservative estimates to evaluate safety bharadhwaj2020conservative.
We also ablate USC by comparing against Uncertain Safety Critic (NR), a USC variant without the uncertainty refinement stage. Other safe exploration approaches discussed in Section 2 were excluded because they represent earlier research from approaches used in our evaluation. To enable a fair comparison, all methods use the same DDPG configuration; algorithm details and hyperparameter information are provided in Appendix D. Following bharadhwaj2020conservative, we perform five independent runs, with results showing mean scores and confidence intervals (standard error).
@@ 5.1. Performance Results
Safety Gymnasium Gymnasium Robotics Mujoco
Technique Metric Goal1 ($\chi=1.0$) Goal2 ($\chi=5.0$) Button1 ($\chi=5.0$) Button2 ($\chi=5.0$) FetchReach ($\chi=5.0$) HalfCheetah ($\chi=0.1$)
DDPG Reward $\uparrow$ $7.90\pm 2.40$ $9.13\pm 2.88$ $5.36\pm 1.87$ $9.09\pm 2.69$ $4.74\pm 0.10$ $9.24\pm 2.17$
Cost $\downarrow$ $3.10\pm 4.11$ $41.80\pm 40.59$ $71.40\pm 126.01$ $31.30\pm 24.58$ $3.09\pm 1.22$ $69.80\pm 27.32$
PID-Lag Reward $\uparrow$ $6.50\pm 2.11$ $8.73\pm 3.51$ $07\pm 0.19 0.0$$07\pm 0.20$ $3.68\pm 3.06$ $5.21\pm 0.47$
Cost $\downarrow$ $4.60\pm 7.00$ $25.10\pm 16.59$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $2.61\pm 1.73$ $0.20\pm 0.40$
CPO Reward $\uparrow$ $6.76\pm 2.52$ $8.04\pm 2.98$ $6.22\pm 2.06$ $7.73\pm 4.52$ $4.80\pm 0.11$ $6.57\pm 0.26$
Cost $\downarrow$ $7.80\pm 11.74$ $21.50\pm 20.41$ $43.70\pm 64.24$ $47.30\pm 84.94$ $2.80\pm 1.17$ $10.90\pm 3.59$
RCPO Reward $\uparrow$ $6.74\pm 2.21$ $8.02\pm 3.76$ $0.01\pm 0.26$ $5.30\pm 12.53$ $4.72\pm 0.16$ $9.48\pm 1.68$
Cost $\downarrow$ $2.40\pm 4.98$ $33.20\pm 53.33$ $0.00\pm 0.00$ $101.80\pm 147.89$ $2.53\pm 1.61$ $0.80\pm 1.17$
SC Reward $\uparrow$ $7.57\pm 1.89$ $11.48\pm 2.51$ $2.63\pm 2.14$ $0.11\pm 0.17$ $4.81\pm 0.09$ $6.91\pm 0.47$
Cost $\downarrow$ $4.80\pm 5.93$ $31.20\pm 33.28$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $3.12\pm 1.09$ $0.50\pm 0.92$
CSC Reward $\uparrow$ $7.21\pm 2.64$ $9.44\pm 2.79$ $3.96\pm 1.62$ $6.33\pm 2.95$ $4.78\pm 0.10$ $8.66\pm 1.57$
Cost $\downarrow$ $2.00\pm 6.00$ $27.00\pm 17.49$ $7.90\pm 15.93$ $14.10\pm 17.39$ $2.98\pm 1.17$ $16.00\pm 10.11$
USC (NR) Reward $\uparrow$ $5.79\pm 2.34$ $9.38\pm 3.95$ $7.65\pm 2.66$ $9.59\pm 2.59$ $4.78\pm 0.13$ $5.74\pm 0.28$
Cost $\downarrow$ $2.60\pm 5.66$ $6.30\pm 26.71 22$$2.20\pm 25.96$ $24.00\pm 41.23$ $2.97\pm 1.75$ $0.60\pm 0.49$
USC Reward $\uparrow$ $7.19\pm 1.97$ $9.48\pm 3.30$ $6.41\pm 2.53$ $8.51\pm 4.79$ $4.78\pm 0.07$ $5.44\pm 0.42$
Cost $\downarrow$ $0.70\pm 2.10$ $20.30\pm 24.63$ $15.90\pm 28.40$ $13.60\pm 13.71$ $2.45\pm 1.25$ $0.20\pm 0.40$
@@
Technique Goal1 Goal2 Button1 Button2 FetchReach
DDPG $\% 100100\%$ $90\%$ $\% 100100\%$
PID-Lag $\% 100100\%$ $\% 00\%$ $90\%$
CPO $\% 100100\%$ $\% 100100\%$ $100\%$
RCPO $\% 100100\%$ $0\%$ $30\%$ $100\%$
SC $\% 100100\%$ $50\%$ $0\%$ $100\%$
CSC $\% 100100\%$ $70\%$ $80\%$ $100\%$
USC (NR) $\% 100100\%$ $\% 100100\%$ $100\%$
USC $\% 100100\%$ $\% 100100\%$ $100\%$
@@ B.1. Theorem 1
Theorem 1. The log-exp-sum (LSE) penalty, used in Equation 3, is more effective at encouraging conservatism when the $m$ alternative actions are sampled uniformly over the action space, rather than from the current policy bharadhwaj2020conservative. If the policy aims to avoid high-cost regions, then the uniform sampling is more likely to draw from high-cost out-of-distribution actions. Formally, for any fixed state $s$, the conservative term is strictly larger under uniform scaling.
@@ C.1. Extended Analysis of Core Results
The deployment-phase evaluations, see in Figure 9, confirm the trends reported in Table 3. Across all CarGoal and CarButton tasks, USC consistently delivers higher reward while keeping costs relatively low compared to all other methods, and it maintains the highest average distance to hazards. Safety Critic again proves brittle, with frequent cost spikes and poor hazard avoidance, while CSC is more stable but overly conservative, often suppressing reward. Notably, in the more difficult CarButton tasks, USC is the only safety-aware method that sustains balanced performance without collapsing into unsafe or overly cautious behaviour and achieves 100% success rate, whereas the other safety-aware methods fail to reach the goal every time. These results reinforce that USC’s uncertainty-modulated training advantage translates into reliable and safe deployment performance.
@@ Appendix D Hyperparameter Analysis and Computational Overheads
Hyperparameters for the DDPG algorithm started with author recommendations ddpg-paper. They were manually tuned afterwards to achieve a high performance on individual environments before tests were carried out, meaning the RL algorithm for all approaches was of high performance and fair comparison.